# VCR: Visual Caption Restoration

### v3

@@ 4.4 Experimental Results
Table 4.4 presents the exact match scores and Jaccard indices for our evaluations. Additionally, Figure 5 illustrates a comparative analysis of our fine-tuned models against their base counterparts across all four VCR settings. We also provide experiment results on smaller test sets of VCR-wiki for faster benchmarking. Please refer to Table A and A in the Appendix.
The VCR task and the VCR-wiki dataset are specifically designed to advance the development of future VLMs. In this section, we conduct a detailed analysis of the performance of current state-of-the-art (SOTA) models on the VCR task, highlighting key insights through comparative evaluations.
@@
Despite observing high-performing models on the $\text{VCR}_{\text{EN, EASY}}$ and $\text{VCR}_{\text{EN, HARD}}$ settings, such as Qwen2-VL, the majority of recent models continue to struggle, particularly under the more demanding hard settings where performance metrics approach zero. This underscores the inherent difficulty of the VCR task and indicates that significant advancements are still required to achieve robust performance across all settings.
@@
A significant performance degradation is observed when models are evaluated on Chinese configurations, despite assertions of basic English-Chinese bilingual capabilities. This decline is particularly surprising given the logographic nature of Chinese characters, which theoretically offer higher recognizability compared to alphabetic scripts (Wu et al., 2024; Zhao et al., 2022). These results indicate a critical need for targeted improvements in multilingual support to ensure consistent performance across different languages.
@@
Model Architecture Influences Performance Gains from VCR-wiki Finetuning.
Finetuning on the VCR-wiki dataset yields varying performance improvements across different model architectures. Specifically: 1) CogVLM2 exhibits substantial performance enhancements across all four settings post-finetuning, suggesting that its architectural design is well-suited for the type of image-text reasoning required by the VCR task when provided with appropriate data and training. 2) MiniCPM-V2.5 shows only marginal performance increases from an already low baseline, particularly in the $\text{VCR}_{\text{ZH, EASY}}$ and $\text{VCR}_{\text{ZH, HARD}}$ settings. This limited improvement indicates potential architectural limitations that hinder its ability to effectively address the complexities of the VCR task. 3) Qwen2-VL-7B maintains relatively high performance both before and after finetuning, implying that the model is sufficiently pre-trained on relevant tasks to perform well on the VCR task without extensive additional training.
@@
Table 2: Performance of vision language models on the VCR task in English (EN) and Chinese (ZH), for easy and hard modes. We label the best result of each setting with bold fonts, and the best open-source model with underline. A superscript of * marks that the model was released after the initial public release of the VCR-wiki dataset (June 10, 2024). Subscripts show bootstrapped standard deviation.
@@
CogVLM2 19B $83.25_{0.07}$ $78.29_{0.04}$ 4.96 $89.75_{0.1}$ $88.07_{0.08}$ 1.68
DeepSeek-VL 1.3B $23.04_{0.05}$ $31.09_{0.12}$8.04 $46.84_{0.07}$ $52.36_{0.06}$5.52
@@
Qwen2-VL* 7B $\underline{89.70_{0.34}}$ $\underline{93.44_{0.26}}$3.74 $\underline{93.84_{0.24}}$ $\underline{97.47_{0.12}}$3.62
@@
CogVLM2 19B $37.98_{0.18}$ ${17.68_{0.06}}$ 20.3 $59.99_{0.05}$ ${39.69_{0.03}}$ 20.3
DeepSeek-VL 1.3B $0.16_{0.01}$ $0.39_{0.02}$0.23 $11.89_{0.02}$ $11.47_{0.03}$ 0.42
@@
Llama-3.2-Vision* 11B $14.09_{0.40}$ $6.92_{0.27}$ 7.17 $35.26_{0.36}$ $26.35_{0.29}$ 8.90
Llama-3.2-Vision* 90B $14.91_{0.40}$ $13.06_{0.37}$ 1.85 $35.44_{0.35}$ $34.44_{0.35}$ 1.00
@@
Qwen2-VL* 7B $\underline{\bm{74.32_{0.47}}}$ $\underline{75.20_{0.49}}$0.88 $\underline{85.47_{0.30}}$ $\underline{87.63_{0.27}}$2.15
@@
Open CogVLM2-Chinese 19B $33.24_{0.04}$ $30.7_{0.07}$ 2.54 $57.57_{0.06}$ $53.66_{0.04}$ 3.91
DeepSeek-VL 1.3B $0.0_{0.0}$ $0.0_{0.0}$ 0 $6.56_{0.01}$ $3.17_{0.02}$ 3.4
@@
Qwen2-VL* 7B $\underline{\bm{59.94_{0.49}}}$ $\underline{\bm{67.48_{0.47}}}$7.54 $\underline{\bm{76.95_{0.32}}}$ $\underline{\bm{82.63_{0.28}}}$5.67
@@
Open CogVLM2-Chinese 19B $1.34_{0.03}$ $2.67_{0.02}$1.32 $17.35_{0.03}$ $19.51_{0.03}$2.16
DeepSeek-VL 1.3B $0.0_{0.0}$ $0.0_{0.0}$ 0 $6.46_{0.01}$ $3.22_{0.02}$ 3.24
@@
Qwen2-VL* 7B $\underline{\bm{18.33_{0.37}}}$ $\underline{\bm{27.58_{0.44}}}$9.26 $\underline{\bm{43.55_{0.34}}}$ $\underline{\bm{54.24_{0.34}}}$10.69
@@ 4.5 Human Evaluation
The human evaluation results are shown in Table 3. Although current SOTA models suffer from the challenge, fluent speakers can easily achieve more than 90 percent accuracy across difficulties. Please refer to Table A to compare all models with human evaluation results using the same test cases.
@@ 5 Related Work
Several datasets have been proposed for visual question answering VQA Antol et al. (2015); Zhang et al. (2016); Goyal et al. (2017); Mishra et al. (2019b). FVQA Wang et al. (2018) and OK-VQAMarino et al. (2019) are datasets about knowledge-based visual question answering and contains questions that necessitate the usage of external knowledge resources. CLEVR Johnson et al. (2017) is a synthetic VQA dataset that mainly focuses on visual reasoning abilities. Recognizing the need to develop VQA models that can understand text, Text-VQA Singh et al. (2019); Biten et al. (2019); Mishra et al. (2019a); Wang et al. (2020) aims to read and reason about texts embedded within images in the context of image-question answering. Several datasets Singh et al. (2019); Biten et al. (2019); Mishra et al. (2019a) have been developed for the Text-VQA task, such as the TextVQA dataset Singh et al. (2019) and the ST-VQA dataset Biten et al. (2019) on natural images, the OCR-VQA dataset Mishra et al. (2019b) on book or movie covers, the InfographicVQA Mathew et al. (2022) dataset on infographics, and the DocVQA dataset Mathew et al. (2021) on document images.
@@
Vision-language models are designed for tasks that involve understanding and generating content from images and text Sun et al. (2023); Liu et al. (2023b); Laurençon et al. (2023; 2024). For example, models have been developed to combine Llama3 with advanced vision-language processing capabilities to handle complex multimodal tasks Yu et al. (2024); Xu et al. (2024); Hu et al. (2023); Yu et al. (2023); Wang et al. (2023b); Dong et al. (2024a). Qwen-VL Bai et al. (2023) enhances visual-linguistic representations for more accurate contextual interpretations, while OpenGVLab-InternVL-Chat Chen et al. (2023; 2024b) merges the InternVL framework with interactive chat capabilities. These studies typically employ a multimodal encoder  (Radford et al., 2021; Zhai et al., 2023; Wu et al., 2022) to process multimodal data, which is then mapped to the same input space of the language model. General-purpose models such as the GPT-4 series models  (Ouyang et al., 2022; OpenAI et al., 2023), the Claude series models  (Anthropic, 2024), the Gemini series models  (Team et al., 2024a) and the Reka series models  (Team et al., 2024b) have also been adapted for vision-language tasks, demonstrating strong performance in multimodal tasks. Finally, DocLLM Wang et al. (2023a) specializes in document understanding by integrating visual and textual data to enhance the interpretation and generation of document-related content. These models collectively represent significant advancements in vision-language integration, contributing unique capabilities and enhancements to the understanding and generation of multimodal information.
@@
OCR Nagy (2000) and its subproblems Howe (2013); Smith (1995); Shafait et al. (2008); Frinken et al. (2011) have been well-studied in the literature in the constrained setting. However, classical OCR methods often cannot perform well on images captured in the wild in an unconstrained setting. Many new methods have been developed for advancing scene-text recognition on camera-captured images Bissacco et al. (2013); Gupta et al. (2016); Huang et al. (2014); Jaderberg et al. (2014); Wang et al. (2012); Shi et al. (2017); Zhou et al. (2017); Lee & Osindero (2016). In addition to the detection and recognition of OCR tasks, visual question answering has emerged as an important downstream task in the OCR literature. With the development of Text-VQA, new methods for improving the reading abilities in VQA utilizing OCR have been proposed. For example, LoRRA Singh et al. (2019) extends a VQA model Pythia Jiang et al. (2018) with an OCR module to better handle Text-VQA tasks. TAP Yang et al. (2021) incorporates scene texts that are generated from OCR engines during pretraining to further improve Text-VQA capabilities.

### v4

@@ 4.4 Experimental Results
For VCR-Hidden, Table 4.4 presents state-of-the-art VLMs’ performance as of February 2025, with more comprehensive results for more models in Table C (English) and Table C (Chinese). In addition, Table 4.4 shows exact match scores and Jaccard indices on VCR-wiki. Figure 5 compares our fine-tuned models against base models across all VCR settings. For faster benchmarking, Tables D and D in the Appendix provide results on smaller VCR-wiki test sets containing 100 or 500 samples. In this section, we analyze models’ performance as of our paper submission date (October 2024) on the VCR task, highlighting key insights through comparative evaluations.
@@
Despite high-performing models like Qwen2-VL excelling in $\text{VCR}_{\text{EN, EASY}}$ and $\text{VCR}_{\text{EN, HARD}}$ settings, most recent models struggle significantly, especially under harder settings where metrics approach zero. This highlights not only the inherent difficulty of the VCR task but also that subpar performance on VCR-wiki stems from a lack of reasoning capabilities or sufficient text-image alignment rather than unfamiliarity with the underlying text, as many VLMs are pretrained on similar data. These results emphasize the need for advancements in VLM designs to achieve robust performance across all settings. Besides, to avoid over-fitting on the train set we released, we also include results from our hidden test set VCR-Hidden, which is not publically accessible. For most recent models, we observe a performance decline on VCR-Hidden compared with VCR-wiki. We will keep updating the both the public VCR-wiki and private VCR-Hidden leaderboards in our Github repository.
@@
A significant performance degradation is observed when models are evaluated on Chinese configurations, despite assertions of basic English-Chinese bilingual capabilities. This decline is particularly surprising given the logographic nature of Chinese characters, which theoretically offer higher recognizability compared to alphabetic scripts (Wu et al., 2024a; Zhao et al., 2022). These results indicate a critical need for targeted improvements in multilingual support to ensure consistent performance across different languages.
@@
Model Design Influences Performance Gains from VCR-wiki Finetuning.
As shown in Figure 5, finetuning on the VCR-wiki dataset yields varying performance improvements across different model designs. Specifically: 1) CogVLM2 demonstrates substantial performance enhancements across all four settings after finetuning, indicating that its overall design may be well-aligned with the image-text reasoning demands of VCR, though further empirical validation is needed to substantiate this hypothesis. 2) MiniCPM-V2.5 shows only marginal performance increases from an already low baseline, particularly in the $\text{VCR}_{\text{ZH, EASY}}$ and $\text{VCR}_{\text{ZH, HARD}}$ settings. This limited improvement indicates potential design limitations that hinder its ability to effectively address the complexities of the VCR task. 3) Qwen2-VL-7B maintains relatively high performance both before and after finetuning, implying that the model is sufficiently pre-trained on relevant tasks to perform well on the VCR task without extensive additional training.
@@
Table 1: Performance of some SOTA-level VLMs (as of February 2025) on the VCR-Hidden in English and Chinese. We label the best result of each setting and metric with bold fonts. A superscript of * marks that the model was released after the initial public release of the VCR-wiki dataset (June 10, 2024). Subscripts show bootstrapped standard deviation. Refer to Table C for the complete results in English and Table C for the complete results in Chinese.
{NiceTabular}
@—c—c—p4.5cmC2cmC2.3cmC2.3cmC1.8cmC2.3cmC2.3cmC1.8cm—@ Language Open/closed source Model name Model size Exact match (%) $\uparrow$ Jaccard index (%) $\uparrow$
$VI$ + $TEI$ $TEI$ $\Delta$ $VI$ + $TEI$ $TEI$ $\Delta$
English Closed Claude 3.7 Sonnet* - $74.00_{2.99}$ $58.50_{3.38}$ 15.50 $81.53_{2.34}$ $61.47_{3.26}$ 20.06
Gemini 2.0 Flash* - $48.50_{3.55}$ $42.00_{3.55}$ 6.50 $66.35_{2.76}$ $61.16_{2.76}$ 5.20
GPT-4o - $41.50_{3.43}$ $45.50_{3.55}$4.00 $48.09_{3.39}$ $51.51_{3.30}$3.43
o1* - $26.50_{3.17}$ $18.50_{2.64}$ 8.00 $31.17_{3.11}$ $22.06_{2.79}$ 9.11
Grok 2 Vision* - $16.50_{2.43}$ $26.00_{3.15}$9.50 $36.05_{2.47}$ $43.83_{2.79}$7.78
Open CogVLM2 19B $53.00_{3.72}$ $51.00_{3.56}$ 2.00 $64.97_{2.76}$ $61.01_{2.98}$ 3.97
DeepSeek-VL2* 28B $23.50_{2.93}$ $32.00_{3.31}$8.50 $31.75_{2.94}$ $45.50_{2.90}$13.75
InternVL2.5* 78B $83.50_{2.70}$ $\bm{79.50_{2.83}}$ 4.00 $91.47_{1.39}$ $\bm{87.60_{1.88}}$ 3.87
Llama-3.2-Vision* 90B $51.00_{3.52}$ $33.50_{3.20}$ 17.50 $63.21_{2.88}$ $50.55_{2.86}$ 12.66
QvQ-Preview* 72B $66.00_{3.40}$ $62.50_{3.25}$ 3.50 $75.76_{2.66}$ $72.57_{2.73}$ 3.19
Qwen2.5-VL* 72B ${\bm{86.50_{2.43}}}$ $\bm{79.50_{2.88}}$ 7.00 ${\bm{92.28_{1.52}}}$ $86.43_{2.01}$ 5.86
Ovis2* 34B $62.00_{3.52}$ $60.50_{3.43}$ 1.50 $73.30_{2.70}$ $71.34_{2.85}$ 1.97
Chinese Closed Claude 3.7 Sonnet* - $3.00_{1.16}$ $0.50_{0.50}$ 2.50 $9.93_{1.44}$ $0.92_{0.52}$ 9.02
Gemini 2.0 Flash* - $1.00_{0.70}$ $1.00_{0.69}$ 0.00 $12.36_{1.11}$ $11.77_{1.09}$ 0.59
GPT-4o - $1.50_{0.89}$ $0.50_{0.50}$ 1.00 $3.78_{1.01}$ $2.22_{0.66}$ 1.56
Open CogVLM2-Chinese 19B $3.00_{1.17}$ $1.50_{0.82}$ 1.50 $16.05_{1.47}$ $13.15_{1.21}$ 2.90
InternVL2.5* 78B $21.00_{2.91}$ $11.50_{2.30}$ 9.50 $46.23_{2.52}$ $34.80_{2.12}$ 11.43
QvQ-Preview* 72B $18.50_{2.75}$ $23.50_{3.01}$5.00 $24.71_{2.73}$ $34.85_{2.90}$10.13
Qwen2.5-VL* 72B $\bm{33.50_{3.35}}$ $\bm{28.00_{3.21}}$ 5.50 $\bm{52.65_{2.78}}$ $\bm{46.83_{2.72}}$ 5.82
Ovis2* 34B $2.00_{1.00}$ $1.00_{0.73}$ 1.00 $14.12_{1.39}$ $13.49_{1.28}$ 0.62
Table 2: Performance of vision language models on VCR-wiki in English and Chinese, for easy and hard modes. We label the best result of each setting and metric with bold fonts, the best open-source model with underline, and the best open-source model released before VCR-wiki’s initial public release (June 10, 2024) with italic font. A superscript of * marks that the model was released after the initial public release of VCR-wiki. Subscripts show bootstrapped standard deviation. For more latest models’ results, please visit our GitHub repository.
@@
CogVLM2 19B $\mathit{83.25_{0.07}}$ $\mathit{78.29_{0.04}}$ 4.96 $\mathit{89.75_{0.1}}$ $\mathit{88.07_{0.08}}$ 1.68
@@
Qwen2-VL* 7B $89.70_{0.34}$ $93.44_{0.26}$3.74 $93.84_{0.24}$ $\underline{97.47_{0.12}}$3.62
Qwen2-VL* 72B $91.30_{0.32}$ $94.64_{0.26}$3.34 $94.04_{0.23}$ $97.42_{0.14}$3.38
Qwen2.5-VL* 7B $\underline{\bm{94.81_{0.25}}}$ $\underline{\bm{93.79_{0.27}}}$ 1.01 $\underline{\bm{98.09_{0.10}}}$ $97.24_{0.13}$ 0.84
Qwen2.5-VL* 72B $91.87_{0.29}$ $87.65_{0.37}$ 4.23 $95.48_{0.18}$ $90.75_{0.30}$ 4.73
@@
CogVLM2 19B $\mathit{37.98_{0.18}}$ ${\mathit{17.68_{0.06}}}$ 20.3 $\mathit{59.99_{0.05}}$ ${\mathit{39.69_{0.03}}}$ 20.3
@@
InternVL-V2* 76B $20.58_{0.39}$ $20.29_{0.39}$ 0.29 $44.59_{0.34}$ $42.86_{0.34}$ 1.73
Llama-3.2* 11B $14.09_{0.40}$ $6.92_{0.27}$ 7.17 $35.26_{0.36}$ $26.35_{0.29}$ 8.90
Llama-3.2* 90B $14.91_{0.40}$ $13.06_{0.37}$ 1.85 $35.44_{0.35}$ $34.44_{0.35}$ 1.00
@@
Qwen2-VL* 7B $74.32_{0.47}$ $\underline{75.20_{0.49}}$0.88 $85.47_{0.30}$ $87.63_{0.27}$2.15
Qwen2-VL* 72B $69.87_{0.52}$ $71.70_{0.49}$1.83 $82.78_{0.33}$ $85.58_{0.28}$2.80
Qwen2.5-VL* 7B $\underline{\bm{80.47_{0.45}}}$ $73.83_{0.48}$ 6.65 $\underline{\bm{91.65_{0.20}}}$ $\underline{87.85_{0.25}}$ 3.80
Qwen2.5-VL* 72B $79.79_{0.45}$ $67.31_{0.54}$ 12.48 $87.91_{0.29}$ $77.12_{0.42}$ 10.78
@@
Open CogVLM2-Chinese 19B $\mathit{33.24_{0.04}}$ $\mathit{30.7_{0.07}}$ 2.54 $\mathit{57.57_{0.06}}$ $\mathit{53.66_{0.04}}$ 3.91
@@
Qwen2-VL* 7B $59.94_{0.49}$ $67.48_{0.47}$7.54 $76.95_{0.32}$ $82.63_{0.28}$5.67
Qwen2-VL* 72B $65.38_{0.46}$ $74.08_{0.44}$8.70 $81.14_{0.28}$ $\underline{\bm{86.78_{0.25}}}$5.64
Qwen2.5-VL* 7B $72.38_{0.46}$ $46.08_{0.49}$ 26.30 $84.66_{0.28}$ $57.03_{0.42}$ 27.63
Qwen2.5-VL* 72B $\underline{\bm{75.82_{0.41}}}$ $\underline{\bm{72.12_{0.43}}}$ 3.70 $\underline{\bm{86.93_{0.25}}}$ $83.40_{0.28}$ 3.54
@@
Open CogVLM2-Chinese 19B $\mathit{1.34_{0.03}}$ $\mathit{2.67_{0.02}}$1.32 $\mathit{17.35_{0.03}}$ $\mathit{19.51_{0.03}}$2.16
@@
Qwen2-VL* 7B $18.33_{0.37}$ $27.58_{0.44}$9.26 $43.55_{0.34}$ $54.24_{0.34}$10.69
Qwen2-VL* 72B $15.30_{0.36}$ $27.35_{0.44}$12.05 $39.71_{0.32}$ $53.63_{0.35}$13.91
Qwen2.5-VL* 7B $19.84_{0.38}$ $12.11_{0.32}$ 7.74 $45.38_{0.35}$ $31.23_{0.34}$ 14.14
Qwen2.5-VL* 72B $\underline{\bm{31.53_{0.47}}}$ $\underline{\bm{29.47_{0.44}}}$ 2.06 $\underline{\bm{56.69_{0.35}}}$ $\underline{\bm{54.41_{0.34}}}$ 2.27
@@ 4.5 Human Evaluation
The human evaluation results are shown in Table 3. Although current SOTA models suffer from the challenge, fluent speakers can easily achieve more than 90 percent accuracy across difficulties. Please refer to Table D to compare all models with human evaluation results using the same test cases.
@@ 5 Related Work
Several datasets have been proposed for visual question answering VQA (Antol et al., 2015; Zhang et al., 2016; Goyal et al., 2017; Mishra et al., 2019b). FVQA (Wang et al., 2018) and OK-VQA(Marino et al., 2019) are datasets about knowledge-based visual question answering and contains questions that necessitate the usage of external knowledge resources. CLEVR (Johnson et al., 2017) is a synthetic VQA dataset that mainly focuses on visual reasoning abilities. Recognizing the need to develop VQA models that can understand text, Text-VQA (Singh et al., 2019; Biten et al., 2019; Mishra et al., 2019a; Wang et al., 2020) aims to read and reason about texts embedded within images in the context of image-question answering. Several datasets (Singh et al., 2019; Biten et al., 2019; Mishra et al., 2019a) have been developed for the Text-VQA task, such as the TextVQA dataset (Singh et al., 2019) and the ST-VQA dataset (Biten et al., 2019) on natural images, the OCR-VQA dataset (Mishra et al., 2019b) on book or movie covers, the InfographicVQA (Mathew et al., 2022) dataset on infographics, and the DocVQA dataset (Mathew et al., 2021) on document images.
@@
Vision-language models are designed for tasks that involve understanding and generating content from images and text (Sun et al., 2023; Liu et al., 2023b; Laurençon et al., 2023; 2024a). For example, models have been developed to combine Llama3 with advanced vision-language processing capabilities to handle complex multimodal tasks (Yu et al., 2024; Xu et al., 2024; Hu et al., 2023; Yu et al., 2023; Wang et al., 2023b; Dong et al., 2024a). Qwen-VL (Bai et al., 2023) enhances visual-linguistic representations for more accurate contextual interpretations, while OpenGVLab-InternVL-Chat (Chen et al., 2023; 2024c) merges the InternVL framework with interactive chat capabilities. These studies typically employ a multimodal encoder (Radford et al., 2021; Zhai et al., 2023; Wu et al., 2022) to process multimodal data, which is then mapped to the same input space of the language model. General-purpose models such as the GPT-4 series models (Ouyang et al., 2022; OpenAI et al., 2023), the Claude series models (Anthropic, 2024), the Gemini series models (Team et al., 2024a) and the Reka series models (Team et al., 2024b) have also been adapted for vision-language tasks, demonstrating strong performance in multimodal tasks. Finally, DocLLM (Wang et al., 2023a) specializes in document understanding by integrating visual and textual data to enhance the interpretation and generation of document-related content. These models collectively represent significant advancements in vision-language integration, contributing unique capabilities and enhancements to the understanding and generation of multimodal information.
@@
OCR (Nagy, 2000) and its subproblems (Howe, 2013; Smith, 1995; Shafait et al., 2008; Frinken et al., 2011) have been well-studied in the literature in the constrained setting. However, classical OCR methods often cannot perform well on images captured in the wild in an unconstrained setting. Many new methods have been developed for advancing scene-text recognition on camera-captured images (Bissacco et al., 2013; Gupta et al., 2016; Huang et al., 2014; Jaderberg et al., 2014; Wang et al., 2012; Shi et al., 2017; Zhou et al., 2017; Lee & Osindero, 2016). In addition to the detection and recognition of OCR tasks, visual question answering has emerged as an important downstream task in the OCR literature. With the development of Text-VQA, new methods for improving the reading abilities in VQA utilizing OCR have been proposed. For example, LoRRA (Singh et al., 2019) extends a VQA model Pythia (Jiang et al., 2018) with an OCR module to better handle Text-VQA tasks. TAP (Yang et al., 2021) incorporates scene texts that are generated from OCR engines during pretraining to further improve Text-VQA capabilities.